Assume you purchase raw materials worth 5 million GHS from a company in Ghana. You have negotiated credit terms that allow you to make payment in 180 days. The spot exchange rate is 1 USD = 5 GHS. However, you are concerned that the currency of Ghana may appreciate in the future and increase your cost of raw materials in dollar terms. You have been unable to find a bank willing to enter into a forward contract with you. There are no options available either. How can you hedge your risk. Interest rates in US and Ghana are 3% and 2%, respectively. The cost of capital for the firm is 9%. Explain your logic in words. Show math. Provide complete answer supported by logic and analysis.
Answer #1

IN the given situation, Hedging can be done by Money Market Hedge.

It is the method of creating an asset or a liability in the currency that is either payable or receivable.

Steps for Money Market Hedge:

Liability after 180 days = 5000000 GHS

Interest Rate for GHS = 2%

Therefore, we should invest 5000000/1.01 = 4950495.05 GHS TODAY to receive 5000000 GHS after 180 days

Today

1) Buy 4950495.05 for 4950495.05/5 = $990099.01

2) Invest GHS 4950495.05 for 180 days

After 180 days

3) Realize the investment proceeds with interest = 4950495.05*1.01 = 5000000 GHS

4) Pay the liability of 5000000 GHS

By undergoing above mentioned hedge, the company can eliminate its FOREIGN EXCHANGE FLUCTUATION RISK. It has locked its exposure at $990099.01 in today's terms and at 990099.01*(1.045) = $1034653.47 after 180 days.